OpenStudy (anonymous):

solve the equation |x+6|-3=13

myininaya (myininaya):

|x+6|=16 x+6=16 , x+6=-16 x=10 , x=-22 check x=10: |10+6|-3=16-3=13 :) check x=-22: |-22+6|-3=|-16|-3=16-3=13 :)

OpenStudy (anonymous):

thank you so much :)